Tsao Company budgets on an annual basis. The following beginning and ending inventory levels (in units) are planned for the year 20x1. Three units of raw material are required to produce each unit of finished product.

Required:

  1. If Tsao Company plans to sell 450,000 units during the year, compute the number of units the firm would have to manufacture during the year.
  2. If 540,000 finished units were to be manufactured by Tsao Company during the year, determine the amount of raw material to be purchased.
January 1 December 31
Raw material 84,000 111,000
Work in process 19,000 19,000
Finished goods 84,000 47,000
Answer #1

1) Compute number of units if company plan to sell 450000units

Goods manufactured = closing balance (finished goods)+unit's sold - opening balance

=47000+450000-84000

= 413000

2) if 540000 finished unit's determine Raw material to be purchased

3units of raw material are required for one unit of finished goods

RAW MATERIAL needed = 540000×3=1620000

Raw material purchase= closing balance of raw material +Raw material consumed - opening balance of Raw material

=111000+1620000 - 84000

=1647000 units

ALL THE BEST

PLEASE DO SUPPORT US

ANY DOUBT PLEASE COMMENT BELOW

THANK YOU